integral(from 4 to 10)([x^2][x^2-28x+196]+[x^2] dx

Grade:10

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
I = \int_{4}^{10}(x^{2}(x^{2}-28x+196) + x^{2})dxI = \int_{4}^{10}x^{2}(x^{2}-28x+197)dx
I = \int_{4}^{10}x^4dx - \int_{4}^{10}28x^{3}dx + \int_{4}^{10}197x^{2}dx
I = (\frac{x^{5}}{5})_{4}^{10} - (7.x^{4})_{4}^{10} + (\frac{197.x^{3}}{3})_{4}^{10}
I = \frac{65256}{5} = 13051.2
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty